# Dark energy under a gauge symmetry: A review of gauged quintessence and its implications

### v1

@@ 2.1 Action and Equations of Motion
which represents how the energy density of quintessence evolves with cosmic expansion. For quintessence to drive cosmic acceleration at late times, one generally requires $w_{\phi}<-\frac{1}{3}$. In the slow-roll limit, where $\dot{\phi}^{2}\ll V_{0}(\phi)$, this parameter approaches $w_{\phi}\simeq-1$, which is the cosmological constant $\Lambda$ limit. Due to the positivity of the kinetic energy, the equation of state for quintessence cannot be less than $-1$, i.e., $w_{\psi}\geq-1$.
@@ 5.1 Review: Misalignment for a Coherent Scalar Field
[…]
@@ 5.2 Extension to a Coherent Vector Field
We now turn to the analogous scenario for a massive spin-1 field $X$[48]. Before inflation, one might imagine that the vector field values are oriented randomly in different patches of the universe. After sufficient inflation, however, the observable universe is dominated by a single direction of $X$ (see Fig. 6). We focus on the zero-momentum mode, treating
@@ 6.1 Evolutionary Stages
[…]
@@ 6.2 Illustrative Numerical Results
The subsequent three panels show the normalized energy densities of $\phi$ and $X$, the field evolution, and the equation of state, respectively. In the later part of Case (ii), $\rho_{X}$ becomes comparable to $\rho_{\phi}$, allowing the gauge potential to affect the evolution of $\phi$. In this regime, $\phi$ may have mild oscillations along the trapped potential, as seen from the third and fourth panels of Fig. 8. As the evolution transitions into Case (iii), where $m_{X}\gtrsim H$, the $X$ field starts oscillation, producing strong oscillation patterns in the $\phi$ field. This leads to a violent energy transfer between $X$ and $\phi$, reflected in the oscillations of $w_{\phi}$. Notably, the value of $w_{\phi}$ can extend beyond the typical range of $[-1,1]$ due to the energy exchange between $\phi$ and $X$.
@@ 7 Hubble Tension
A persistent discrepancy in measurements of the Hubble constant $H_{0}$ has emerged in recent years, dubbed the “Hubble tension.” Specifically, early-Universe probes (such as fits to the cosmic microwave background (CMB) data under $\Lambda$CDM) favor $H_{0}\approx 67\,\mathrm{km\,s^{-1}\,Mpc^{-1}}$[37], whereas direct late-Universe measurements (e.g., using standard candles in the local universe) indicate $H_{0}\approx 73\,\mathrm{km\,s^{-1}\,Mpc^{-1}}$[49]. The mismatch between these two primary methods suggests new physics beyond the standard $\Lambda$CDM model or unrecognized systematic errors in one (or both) measurements.
@@ 7.1 BAO Constraints and Angular Sound Horizon
[…]
@@ 7.2 Why $w<-1$ at Late Times?
and $\rho\propto a^{-3(1+w)}$, it follows that a higher present-day $H_{0}$ can be balanced by a period in which $\rho_{\mathrm{DE}}$ evolves more slowly or even increases with decreasing redshift. This generally points to an effective equation of state $w_{\mathrm{eff}}<-1$ for part of the late-time evolution. Unfortunately, uncoupled quintessence has $w\geq-1$, making the tension worse compared to $\Lambda$CDM.[56, 55]
@@ 7.3 Effective Equation of State in Gauged Quintessence
Several authors have explored interacting dark energy scenarios in which $w_{\mathrm{eff}}$ can dip below $-1$, potentially alleviating the Hubble tension. In our gauged quintessence framework, the dynamical mass of the dark gauge boson $m_{X}^{2}(t)=g_{X}^{2}\,\phi^{2}(t)$ induces an energy flow between $\phi$ and $X$. As a result, the effective dark energy density $\rho_{\widetilde{\mathrm{DE}}}$ can behave differently from standard quintessence.
Figure 10 provides a schematic illustration of how $H(z)$ might differ from $\Lambda$CDM (black curves) if an extra interaction shifts energy between $\phi$ and other components. Meanwhile, Fig. 11 shows the effective equation of state parameter $w_{\mathrm{eff}}(\widetilde{\mathrm{DE}})$ for gauged quintessence under various assumptions about the fraction of $X$ in the total dark matter. A convenient expression is
@@
where $w_{0}=-1$ in $\Lambda$CDM, $m_{X}^{0}$ and $\rho_{X}^{0}$ are the present-day mass and density of $X$, and $\rho_{\widetilde{\mathrm{DE}}}$ includes both $\phi$ and the portion of $X$ that effectively contributes to late-time acceleration. Because $m_{X}$ was smaller in the past ($m_{X}<m_{X}^{0}$), the overall correction in Eq. (27) can be negative, potentially driving $w_{\mathrm{eff}}$ below $-1$.
Refer to caption($\Lambda$CDM)(Larger $H_{0}$)
@@
Figure 10: Conceptual sketch of the expansion rate $H(a)$ in $\Lambda$CDM (black curves) and in a hypothetical dark energy model that reduces $H(a)$ at intermediate redshifts, then yields a higher $H_{0}$. Such a scenario may mitigate the Hubble tension but must remain consistent with the BAO scale.
@@ 7.4 Numerical Indications and Outlook
In Fig. 11, we show that for a sufficiently large fraction of $X$ within the dark matter component, $w_{\mathrm{eff}}\bigl{(}\widetilde{\mathrm{DE}}\bigr{)}$ can indeed dip below $-1$. This behavior is absent in uncoupled quintessence and may help reconcile the discrepancy between early and late measurements of $H_{0}$. A detailed resolution of the Hubble tension, however, would require a full numerical fit to CMB data and other large-scale structure constraints, which we reserve for future work.
@@
Figure 11: The effective equation of state $w_{\mathrm{eff}}\bigl{(}\widetilde{\mathrm{DE}}\bigr{)}$ for gauged quintessence.[13] Different colored lines denote varying fractions of $X$ in the total dark matter content. Values dipping below $-1$ in recent cosmological time ($z\lesssim 1$) may help ease the Hubble tension.
In summary, gauged quintessence could potentially serve as a concrete example of the interacting dark energy model that addresses the Hubble tension. By allowing energy transfer between $\phi$ and a mass-varying vector boson $X$, the effective equation of state for dark energy can be driven below $-1$ for a period, lowering $H(z)$ at intermediate redshifts while retaining a larger final value $H_{0}$. Verifying this conjecture demands a systematic comparison with precision cosmological data, including BAO and CMB measurements.
@@ 8.4 Constraints and Future Prospects
Figure 14: A representative overview of constraints on the kinetic mixing parameter $\varepsilon$ and the dark gauge coupling $g_{X}$.[15] The orange region is excluded by current X-ray/$\gamma$-ray diffuse background data. The blank region remains open to discovery, offering interesting prospects for future surveys.
In summary, gauged quintessence models allow the dark gauge boson $X$ to carry a time-dependent mass. Once a kinetic mixing portal is introduced, these mass-varying dark photons can be produced or decayed at different epochs with shifting kinematics, thus imprinting unique signatures in various electromagnetic channels. Ongoing and next-generation observations of diffuse photon backgrounds, as well as high-precision $X$-ray/$\gamma$-ray instruments, can probe sizable portions of the $(\varepsilon,g_{X})$ parameter space, opening a new window into the physics of dark energy beyond purely gravitational effects.
@@ 9 Concluding Remarks
Much like the ongoing, extensive efforts to uncover the nature of dark matter, the search for a governing symmetry in the dark energy sector remains open-ended. As this line of inquiry continues, more symmetries may come to light, and a fully unified model of the universe could ultimately integrate these new insights. We hope that gauged quintessence will serve as a useful paradigm for exploring physics beyond the minimal $\Lambda$CDM scenario, motivating future work in both theoretical and observational domains.
Acknowledgments. The material in this review is based on Refs. 13, 14, 15. This work was supported in part by the National Research Foundation of Korea (Grant No. RS-2024-00352537).

### v2

@@ 2.1 Action and Equations of Motion
which represents how the energy density of quintessence evolves with cosmic expansion. For quintessence to drive cosmic acceleration at late times, one generally requires $w_{\phi}<-\frac{1}{3}$. In the slow-roll limit, where $\dot{\phi}^{2}\ll V_{0}(\phi)$, this parameter approaches $w_{\phi}\simeq-1$, which is the cosmological constant $\Lambda$ limit. Due to the positivity of the kinetic energy, the equation of state for quintessence cannot be less than 1-1, i.e., $w_{\psi}\geq-1$.
@@ 5.1 Review: Misalignment for a Coherent Scalar Field
$\ddot{\varphi}+3H\dot{\varphi}+m_{\varphi}^{2}\,\varphi\;=\;0,\qquad\rho_{\varphi}\;=\;\frac{1}{2}\,\dot{\varphi}^{2}+\frac{1}{2}\,m_{\varphi}^{2}\,\varphi^{2},$ (19)
@@ 5.2 Extension to a Coherent Vector Field
We now turn to the analogous scenario for a massive spin-1 field XX[48]. Before inflation, one might imagine that the vector field values are oriented randomly in different patches of the universe. After sufficient inflation, however, the observable universe is dominated by a single direction of $X$ (see Fig. 6). We focus on the zero-momentum mode, treating
@@ 6.1 Evolutionary Stages
    $X_{i}(t)\propto\text{Re}\left[\frac{\chi_{i}}{\sqrt{am_{X}}}e^{i\int dt\,m_{X}}\right].$ (25)
@@ 6.2 Illustrative Numerical Results
The subsequent three panels show the normalized energy densities of $\phi$ and $X$, the field evolution, and the equation of state, respectively. In the later part of Case (ii), $\rho_{X}$ becomes comparable to $\rho_{\phi}$, allowing the gauge potential to affect the evolution of $\phi$. In this regime, $\phi$ may have mild oscillations along the trapped potential, as seen from the third and fourth panels of Fig. 8. As the evolution transitions into Case (iii), where $m_{X}\gtrsim H$, the $X$ field starts oscillation, producing strong oscillation patterns in the $\phi$ field. This leads to a violent energy transfer between $X$ and $\phi$, reflected in the oscillations of $w_{\phi}$. Notably, the value of $w_{\phi}$ can extend beyond the typical range of [1,1][-1,1] due to the energy exchange between $\phi$ and $X$.
@@ 7 Hubble Tension
[…]
@@ 7.1 BAO Constraints and Angular Sound Horizon
$\begin{split}r_{s}&=\int_{z_{s}}^{\infty}\frac{c_{s}(z)}{H(z)}\,dz,\\ D(z)&=\int_{0}^{z}\frac{c}{H(z)}\,dz\;=\;\frac{r_{s}}{\theta},\end{split}$ (26)
@@ 7.2 Why w<1w<-1 at Late Times?
[…]
@@ 7.3 Effective Equation of State in Gauged Quintessence
Several authors have explored interacting dark energy scenarios in which $w_{\mathrm{eff}}$ can dip below 1-1, potentially alleviating the Hubble tension. In our gauged quintessence framework, the dynamical mass of the dark gauge boson $m_{X}^{2}(t)=g_{X}^{2}\,\phi^{2}(t)$ induces an energy flow between $\phi$ and $X$. As a result, the effective dark energy density $\rho_{\widetilde{\mathrm{DE}}}$ can behave differently from standard quintessence.
[…]
@@
[…]
@@ 7.4 Numerical Indications and Outlook
In Fig. 11, we show that for a sufficiently large fraction of $X$ within the dark matter component, $w_{\mathrm{eff}}\bigl{(}\widetilde{\mathrm{DE}}\bigr{)}$ can indeed dip below 1-1. This behavior is absent in uncoupled quintessence and may help reconcile the discrepancy between early and late measurements of $H_{0}$. A detailed resolution of the Hubble tension, however, would require a full numerical fit to CMB data and other large-scale structure constraints, which we reserve for future work.
@@
Figure 11: The effective equation of state $w_{\mathrm{eff}}\bigl{(}\widetilde{\mathrm{DE}}\bigr{)}$ for gauged quintessence.[13] Different colored lines denote varying fractions of $X$ in the total dark matter content. Values dipping below 1-1 in recent cosmological time ($z\lesssim 1$) may help ease the Hubble tension.
In summary, gauged quintessence could potentially serve as a concrete example of the interacting dark energy model that addresses the Hubble tension. By allowing energy transfer between $\phi$ and a mass-varying vector boson $X$, the effective equation of state for dark energy can be driven below 1-1 for a period, lowering $H(z)$ at intermediate redshifts while retaining a larger final value $H_{0}$. Verifying this conjecture demands a systematic comparison with precision cosmological data, including BAO and CMB measurements.
@@ 8.4 Constraints and Future Prospects
Figure 14: A representative overview of constraints on the kinetic mixing parameter $\varepsilon$ and the dark gauge coupling $g_{X}$.[15] The orange region is excluded by current X-$\gamma-ray$ diffuse background data. The blank region remains open to discovery, offering interesting prospects for future surveys.
In summary, gauged quintessence models allow the dark gauge boson $X$ to carry a time-dependent mass. Once a kinetic mixing portal is introduced, these mass-varying dark photons can be produced or decayed at different epochs with shifting kinematics, thus imprinting unique signatures in various electromagnetic channels. Ongoing and next-generation observations of diffuse photon backgrounds, as well as high-precision $X$-$\gamma-ray$ instruments, can probe sizable portions of the $(\varepsilon,g_{X})$ parameter space, opening a new window into the physics of dark energy beyond purely gravitational effects.
@@ 9 Concluding Remarks
[…]
Acknowledgments. The material in this review is based on Refs. [13, 14, 15]. This work was supported in part by the National Research Foundation of Korea (Grant No. RS-2024-00352537).